substitution
a - 4b = 2
2a - 5b = 1
with steps too

Answer by mananth(16946)

1a+4b=2 .............1
-5b=1-2a ---------2
/-5=
b=-0.2+0.4a
Plug the value of y in (1)
1a+4(-0.2+0.4a)=2
1a-0.8+1.6a=2
1a+1.6a=2+0.8
2.6a=2.8
/2.6
a=1.08
Plug the value of x in (1)
1a+4b=2
1*1.08+4b=2
1.08+4b=2
4b=0.92
/4
b=0.23